You are to drive 300 km to an interview. The interview is (e its at 11:15 A.M. You plan to drive at 100 km/h, so you leave at 8:00 d A.M. to allow some extra time. You drive at that speed for the first 100 km, but then construction work forces you to slow to 40 km/h for 40 km. What would be the least speed needed for the rest of the trip to arrive in time for the interview?
Answer #1

Total time required to travel 300 km is

T=11.15-8 =3 hours 15 minutes

in hours

T=3+(15/60) =3.25 hours

Distance covered in first hour

d1=V1t1 = 100*1 =100 km

Distance covered in second hour

d2=V2t2 =40*1 =40 km

Remaining time

t3=3.25-1-1=1.25 hours

Remaining distance to be covered

d3=300-100-40 =160 km

Minimum speed required to reach interview in time is

V3=d3/t3=160//1.25 =128 km/h